Operations Report – March 17th

Operations Report 17-MAR-2020

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Monday Generator

6 am SOC was 77

10 am SOC was 73

1 PM SOC was 100

7 pm SOC was 98

Generator will not be run overnight.

Tuesday Generator

8 am SOC was 70

5 pm SOC was 100

Generator was taken back to Grand Junction today and we are on solar

power only.

Diesel Reading – 66%

Station Propane Reading – 69 %

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): 20 gallons

Water Meter: n/a units

Water (static tank) – 250 gallons

Static to Loft Pump used – no

Water in GreenHab – basically empty gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: no

Hours: 152.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: no

Hours: 104.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Spirit rover: no

Hours: 134.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: no

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: no

Hours: 139.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: No

Reason for use: n/a

Oil Added? no

ATV Fuel Used: None

Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report.

HabCar used and why, where? Not used

CrewCar used and why, where? CrewCar is currently in Hanksville, we

retrieved it from Grand Junction today.

General notes and comments: Nothing to report

Summary of internet: Nominal

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Plants have been donated to the

Hanksville Community Garden. We will start removing them tomorrow.

Heater has been turned off.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Happy St.

Patrick’s Day from the Red Planet!